# SingingHead: A Large-scale 4D Dataset for Singing Head Animation

### v2

@@ Abstract
Singing, as a common facial movement second only to talking, can be regarded as a universal language across ethnicities and cultures, plays an important role in emotional communication, art, and entertainment. However, it is often overlooked in the field of audio-driven facial animation due to the lack of singing head datasets and the domain gap between singing and talking in rhythm and amplitude. To this end, we collect a high-quality large-scale singing head dataset, SingingHead, which consists of more than 27 hours of synchronized singing video, 3D facial motion, singing audio, and background music from 76 individuals and 8 types of music. Along with the SingingHead dataset, we argue that 3D and 2D facial animation tasks can be solved together, and propose a unified singing facial animation framework named UniSinger to achieve both singing audio-driven 3D singing head animation and 2D singing portrait video synthesis. Extensive comparative experiments with both SOTA 3D facial animation and 2D portrait animation methods demonstrate the necessity of singing-specific datasets in singing head animation tasks and the promising performance of our unified facial animation framework. The dataset and codes will be released for research purposes at: https://wsj-sjtu.github.io/SingingHead/.
Figure 1: We present a new facial animation dataset, SingingHead, which contains more than 27 hours’ synchronized singing video, 3D facial motion, singing audio, and background music (BGM) collected from 76 subjects. Along with the SingingHead dataset, we propose a unified framework, UniSinger, to generate both 3D facial motion and 2D singing portrait video according to the input singing audio.
1 Introduction
@@
Audio-driven facial animation is a challenging task with numerous applications in virtual avatars, extended reality (XR), entertainments, etc. Recently, impressive progress has been made in speech-driven talking head animation [58, 30, 12, 53] thanks to the release of a variety of talking head datasets [36, 59, 50, 13, 8]. Singing, as another common facial movement, plays a significant role in art and entertainment. People like to express themselves or celebrate through singing on various occasions such as social media and gala, which can be seen as a universal language across ethnicities and cultures. However, how to animate a virtual avatar to sing has been rarely explored.
@@
Singing is quite different from speaking in rhythm, expression and amplitude. Therefore, existing methods trained on talking head datasets encounter difficulties when directly applied to the singing task. To solve this problem, it is necessary to collect a singing-specific dataset. Almost all previous works related to audio-driven facial animation focus on talking, and can be divided into two branches: (1) 2D portrait animation, which aims at synthesizing a realistic talking video from a reference portrait image and a driven audio [40, 62, 60, 56, 44, 53, 30]; (2) 3D facial animation, i.e. generate 3D facial motion given speech audio [8, 42, 12, 53, 38]. Only a few works focused on singing [20, 28], however, the datasets they presented are either small in size or of low quality, and none of them provides 2D portrait singing videos (see Tab.1), which seriously limits the performance of singing-audio-driven facial animation and the use of these datasets. The aforementioned limitations severely hinder the development of singing head animation and the flourishing of the automatic virtual avatar singing industry for entertainment.
To solve the problem of the scarcity of singing head datasets and few studies on singing head animation tasks, we present SingingHead, a lab-collected large-scale singing head animation dataset, which contains synchronized singing video, 3D facial motion, singing audio, and background music. Concretely, the design of our dataset considers the following two aspects. (1) Data quality. we recruited volunteers to collect singing videos in our laboratory rather than collect in-the-wild videos from the internet to ensure the acquisition of synchronized high-resolution ($3840\times 2160$) singing video and dry singing audio, together with an accurate 3D face scan of each volunteer for accurate 3D shape estimation. The captured videos are cropped out the facial region and serialized to portrait videos with a resolution of $1024\times 1024$, which form the 2D video part of our SingingHead dataset. Moreover, accurate 3D facial motion represented by FLAME [27] parameters are estimated from the portrait videos with the help of the 3D face scan of each individual. (2) Data diversity. To ensure the diversity of our dataset, we collected singing data from as many as 76 individuals including professional singers and amateur singers. We carefully picked out 8 common music classes, i.e. art music, country music, hip-hop, rock music, folk music, rhythm and blues, theater and drama, and march. We asked the volunteers to choose songs they are familiar with in these classes. In summary, our SingingHead dataset is the first high-quality dataset that contains both 2D portrait videos and 3D facial motion for singing head animation, which is also comparable in size to the latest 2D and 3D facial animation datasets, as illustrated in Tab. 1.
Along with the SingingHead dataset, we propose a unified framework named UniSinger to solve both the 3D singing head animation and 2D singing portrait animation tasks, as shown in Fig. 1. For 3D singing head animation, we propose a transformer-based VAE model [49, 25, 39] to generate diverse 3D facial motion according to the input audio and FLAME [27] shape parameter. For 2D singing portrait animation, given a single portrait image and singing audio, UniSinger can synthesize singing videos of the person in the reference image which is synchronized with the input. Specifically, we first estimate the shape and camera parameters of the input image. Then the input audio and estimated shape parameter are fed to the VAE-based 3D facial animation module to generate 3D facial motion represented by FLAME expression and pose parameters, which is further rendered to depth video using the estimated camera parameter. The depth video together with the input image are fed into the face rendering module to generate the final singing portrait video.
To summarize, the main contributions of this paper are:
  • A large-scale singing head animation dataset, which contains 76 subjects’ more than 27 hours’ synchronized singing video, 3D facial motion, singing audio, and background music.
  • A unified framework, UniSinger, to solve both 3D and 2D singing head animation tasks, i.e., singing audio-driven 3D facial animation and singing audio-driven single portrait image animation.
  • The proposed framework achieves superior performance on 3D singing facial motion generation and 2D singing portrait video synthesizing compared with SOTA methods.
@@
Table 1: Comparison of Talking/Singing Datasets. “Subj.”, “BGM”, “2D”, and “3D” represent subjects, background music, 2D videos, and 3D facial motions, respectively.
Figure 2: Data collection pipeline and data statistics. (a) We use microphone, Azure Kinect camera and light stage to collect raw data including singing audio, video and 3D scan. Then, we process the raw data to obtain the cleaned singing audio, portrait video and 3D facial motion, which make up our SingingHead dataset. (b) We show the pie charts of the song category, the professional distribution of volunteers, and the language, respectively.
2 Related Work
2.1 Audio-driven 3D Facial Animation
In the past few years, audio-driven 3D facial animation [5, 48, 23, 61] has attracted increasing popularity which can be used to animate various 3D face models [3, wu2023ganhead, 18]. According to the input audio, researchers attempt to obtain animation results with realistic expressions and synchronized mouth movements. Benefiting from VOCASET Datset, VOCA [8] proposes a speaker-dependent facial animation pipeline that enables generalization across subjects. FaceFormer [12] introduces transformer [49] to capture long-term audio context and achieve temporary consistent animation performance. However, both of these two works lack attention to upper-face motions. To alleviate this problem, Meshtalk [42] learns a categorical latent space via a dedicated cross-modality loss. To further enhance the realism, Codetalker [53] utilizes a discrete codebook method which successfully avoids over-smoothed animation results, while Emotalk [38] focuses on the disentanglement of emotion and content of the audio.
@@ 2.1 Audio-driven 3D Facial Animation
2.2 Audio-driven 2D Portrait Animation
@@
In recent years, there has been a growing interest in 2D portrait animation task [46, 26, 11, 63] which aims to generate lifelike videos synchronized with the provided audio. Through the integration of a pre-trained Lip-Sync expert [7], Wav2Lip [40] enhances synchronization accuracy of the generated results. Then MekeIttalk [62] improves the realism via a speaker-aware audio encoder that enables personalized head pose movement. To enable natural head pose movement, Zhou et al. [60] propose a low-dimensional pose code to disentangle and control head poses. With the development in generation and reconstruction tasks, many methods [56, 44, 16, 55, 29, 47] introduce Neural Radiance fields [35] or diffusion model [17] to refine the quality of the generated videos. However, these methods typically suffer from less efficient generation processes due to the time-consuming volume rendering or diffusion sampling procedures. Moda [30] proposes a dual-attention module to estimate dense facial landmarks of deterministic and probabilistic movements to generate natural and detailed results.
All these methods mentioned above mainly focus on speech audio, while other common types of audio such as songs still lack exploration. Since singing audio has a different rhythm than speech audio, and the expressions when singing are more extreme, it is challenging for current methods to directly generalize to the singing head generation task. Therefore, we propose a large singing head dataset to enable diverse types of audio-driven generation.
2.3 2D and 3D Talking Head Datasets
@@
The talking head datasets can be divided into 2D and 3D domains with numerous applications on face modeling and animation [22, 41, 54, 34, 32]. 2D datasets [31, 50, 40, 36, 4, 21, 59] focus on collecting more diverse and higher quality photorealistic data. LRW dataset [40] contains 173 hours of low-resolution videos. To enlarge the data scale, Voxceleb dataset [36] collected more than 2k hours of videos from YouTube. In addition, several works [4, 31, 50] explore collecting emotional talking head data with emotion labels. The SAVEE [21] dataset is the first emotional dataset with only 4 actors. Mead [50] dataset builds a high-resolution dataset with various video clips and viewpoints, covering a large number of large emotional audio-visual corpus. Numerous datasets [8, 42, 13] have also been proposed for 3D talking head, MeshTalk dataset [42] contains detailed meshes topology and proposes a convolution neural network based method. Vocaset dataset [8] constructs dataset by registering meshes to the FLAME [27] topology. Recently, 3D-ETF dataset is proposed by collecting emotional 3D data. Considering numerous datasets, however, almost all the datasets are collected for general talking, there are only a few datasets for 2D/3D singing heads. MusicFace [28] constructs a 2D dataset with only 6 singers by collecting the audio and videos separately. Song2Face dataset [20] collects a 2-hour dataset with background music, and 3D blend shapes from 7 singers. Different from them, our SingingHead dataset is the first singing head dataset containing both 2D videos and accurate 3D face motions with the largest scale (27 hours) and diversity (76 singers).
3 SingingHead Dataset
Since there is no publicly available face animation dataset focus on singing, we devise a data collection pipeline and collect a singing head dataset, as illustrated in Fig. 2.
3.1 Data Acquisition
@@
Design policy. We design the collection of SingingHead dataset by considering the following three points. (1) Variety of songs. For the diversity of the singing dataset, we carefully selected 8 common music categories, i.e., art music, country music, hip-hop, rock music, folk music, rhythm and blues, theater and drama, march, and asked the volunteers to choose 5 to 10 songs they were familiar with and fall to these categories. The pie chart at the top of Fig. 2 (b) illustrates the categories and distributions of the songs we have finally collected. (2) Professionalism of volunteers. Different from talking, singing is a form of art that requires practice, which means that there is a great difference in singing skills between those who practice singing frequently and those who sing less. To this end, we select 50 professional singers and 26 amateurs as our volunteers, and recorded a total of 265 songs performed by professional volunteers and 182 songs performed by amateur volunteers, as illustrated at the bottom of Fig. 2 (b). (3) Acquisition of dry sound. In normal circumstances, singers need to listen to the accompaniment or the original song when singing. In order to record the dry sound of singing, we equipped the volunteers with small Bluetooth earphones to play the accompaniment, and then recorded their voices via another microphone.
@@
3.2 Data Processing
@@
The raw data includes RGB videos, singing dry sounds, accompaniments, and 3D scans. We processed the raw data as illustrated in Fig. 2 (a) to get our final dataset SinginHead, which consists of singing audio, portrait video, and 3D facial motion sequence.
@@
Video cropping and serialization. Since the video we collected has a wide field of view, including the entire upper body, and the volunteers routinely wiggled as they recorded the song, it is necessary to crop out the facial region and serialize the long recording to ensure the face remains in a stable bounding box. To this end, we first detect the bounding box of the facial region using MTCNN [57], then leave a margin around the bounding box to determine the crop box and crop the original video into square portrait video, finally the portrait video is resized to $1024\times 1024$. During the cropping process, we use Dlib [24] to detect the 68 landmarks every 6 frames, and calculate whether the landmarks on the edge are too close to the bounding box, if so, we cut the video at this frame, and start a new video sequence with a newly detected bounding box.
@@
3D facial motion reconstruction. We utilize the cropped portrait videos and the corresponding 3D scans to obtain accurate 3D singing facial motion sequences. Specifically, the 3D facial motion is represented using the expression and pose parameters of the commonly used parametric head model FLAME [27]. To estimate accurate FLAME parameters of the singing head, we first estimate the shape, expression, pose, and camera parameters separately for each frame in the portrait video with a SOTA monocular 3D face reconstruction method EMOCA v2 [9, 14, 15], and average the estimated shape parameters of all frames of a person as his/her shape parameter. Considering that estimating 3D shape from the 2D image is an ill-posed problem, and it is hard to obtain very accurate results, we further use the collected fine-gained 3D scans to fit a more accurate FLAME model by minimizing the 3D landmark loss and mesh-to-scan distance with the shape parameter initialized as the monocular estimated shape. Finally, we use the fitted FLAME shape as the final shape parameter, and reproject the estimated FLAME head back to the original portrait video to optimize the expression, pose, and camera parameters via landmark reprojection loss and smoothing loss (see Sup. Mat. for details). After these steps, we can get the highly accurate 3D singing facial motion represented by FLAME [27] parameters.

### v3

@@ Abstract
Singing, as a common facial movement second only to talking, can be regarded as a universal language across ethnicities and cultures, plays an important role in emotional communication, art, and entertainment. However, it is often overlooked in the field of audio-driven facial animation due to the lack of singing head datasets and the domain gap between singing and talking in rhythm and amplitude. To this end, we collect a high-quality large-scale singing head dataset, SingingHead, which consists of more than 27 hours of synchronized singing video, 3D facial motion, singing audio, and background music from 76 individuals and 8 types of music. Along with the SingingHead dataset, we benchmark existing audio-driven 3D facial animation methods and 2D talking head methods on the singing task. Furthermore, we argue that 3D and 2D facial animation tasks can be solved together, and propose a unified singing head animation framework named UniSinger to achieve both singing audio-driven 3D singing head animation and 2D singing portrait video synthesis, which achieves competitive results on both 3D and 2D benchmarks. Extensive experiments demonstrate the significance of the proposed singing-specific dataset in promoting the development of singing head animation tasks, as well as the promising performance of our unified facial animation framework. The dataset will be released for research purposes at: https://wsj-sjtu.github.io/SingingHead/.
Index Terms:
Singing head, audio-driven 3D facial animation, audio-driven portrait video synthesis, dataset and benchmark.
I Introduction
@@
Audio-driven facial animation is a challenging task with numerous applications in virtual avatars, extended reality (XR), entertainments, etc. Recently, impressive progress has been made in speech-driven talking head animation [1, 2, 3, 4, 5, 6] thanks to the release of a variety of talking head datasets [7, 8, 9, 10, 11]. Singing, as another common facial movement, plays a significant role in art and entertainment. People like to express themselves or celebrate through singing on various occasions such as social media and gala, which can be seen as a universal language across ethnicities and cultures. However, how to animate a virtual avatar to sing has been rarely explored.
@@
Singing is quite different from speaking in rhythm, expression, and amplitude (see Fig. 2). Therefore, existing methods [4, 5, 12, 13, 14, 15, 16, 1, 17] trained on talking head datasets encounter difficulties when directly applied to the singing task. To solve this problem, it is necessary to collect a singing-specific dataset. Almost all previous works related to audio-driven facial animation focus on talking, and can be divided into two branches: (1) 3D facial animation, i.e. generating 3D facial motion given speech audio [11, 18, 4, 5, 19, 12, 13, 14, 20, 21, 22]; (2) 2D portrait animation, which aims at synthesizing a realistic talking video from a reference portrait image and a driven audio [23, 24, 25, 26, 27, 28, 29, 5, 2, 17, 3, 30, 31]. Only a few works focus on singing [32, 33], however, the datasets they presented are either small in size or of low quality, and none of them provide both 2D portrait videos and 3D facial motions (see Tab.I), which seriously limits the performance of singing-audio-driven facial animation and the use of these datasets. The aforementioned limitations severely hinder the development of singing head animation and the flourishing of the automatic virtual avatar singing industry for entertainment.
TABLE I: Comparison of Talking/Singing Datasets. “Subj.”, “BGM”, “2D”, and “3D” represent subjects, background music, 2D videos, and 3D facial motions, respectively.
@@
To solve the problem of the scarcity of singing head datasets and few studies on singing head animation tasks, we present SingingHead, a lab-collected large-scale singing head animation dataset, which contains synchronized singing video, 3D facial motion, singing audio, and background music. Concretely, the design of our dataset considers the following two aspects. (1) Data quality. We recruit volunteers to collect singing videos in our laboratory rather than collecting in-the-wild videos from the internet to ensure the acquisition of synchronized high-resolution ($3840\times 2160$) singing video and dry singing audio, together with an accurate 3D face scan of each volunteer for accurate 3D shape estimation. The captured videos are cropped to obtain the facial region and serialized into portrait videos with a resolution of $1024\times 1024$, which form the 2D video part of our SingingHead dataset. Moreover, accurate 3D facial motion represented by FLAME [35] parameters are estimated from the portrait videos with the help of the 3D face scan of each individual. (2) Data diversity. To ensure the diversity of our dataset, we collect singing data from as many as 76 individuals including professional singers and amateur singers. We carefully pick out 8 common music classes, i.e. art music, country music, hip-hop, rock music, folk music, rhythm and blues, theater and drama, and march, and ask the volunteers to choose songs they are familiar with in these classes. In summary, our SingingHead dataset is the first high-quality dataset that contains both 2D portrait videos and 3D facial motions for singing head animation, which is also comparable in size to the latest 2D and 3D facial animation datasets, as illustrated in Tab. I.
Figure 1: We present a new facial animation dataset, SingingHead, which contains more than 27 hours of synchronized singing video, 3D facial motion, singing audio, and background music (BGM) collected from 76 subjects. Along with the SingingHead dataset, we propose a unified framework, UniSinger, to generate both 3D facial motion and 2D singing portrait video according to the input singing audio.
Along with the SingingHead dataset, we benchmark existing audio-driven 3D facial animation methods and 2D talking head methods on the singing task, which shows less satisfactory performance. To this end, we further propose a unified framework named UniSinger as a baseline method to solve both 3D singing head animation and 2D singing portrait video synthesis tasks, as shown in Fig. 1. For 3D singing head animation, we propose a transformer-based VAE model [36, 37, 38] to generate diverse 3D facial motion according to the singing audio, background music, and FLAME [35] shape parameter. For 2D singing portrait video synthesis, given a single portrait image and input audio, UniSinger can synthesize singing videos of the person in the reference image which are synchronized with the input audio. Specifically, we first estimate the shape and camera parameters of the input image. Then the singing audio, background music, and estimated shape parameter are fed into the VAE-based 3D facial animation module to generate 3D facial motion represented by FLAME expression and pose parameters, which is further rendered to depth video using the estimated camera parameters. The depth video together with the input reference image is fed into the face renderer module to generate the final singing portrait video.
To summarize, the main contributions of this paper are:
  • A large-scale singing head animation dataset, SingingHead, which contains 76 subjects’ more than 27 hours of synchronized singing video, 3D facial motion, singing audio, and background music.
@@
  • Two benchmarks for singing head animation: singing audio-driven 3D singing head animation and 2D singing portrait video synthesis.
  • A unified framework, UniSinger, which is able to solve both 3D and 2D singing head animation tasks and achieves competitive results on both benchmarks.
II Related Work
II-A Audio-driven 3D Facial Animation
In the past few years, audio-driven 3D facial animation [39, 40, 41, 42] has attracted increasing popularity which can be used to animate various 3D head models [43, 44, 45, 46, 47]. According to the input audio, researchers attempt to obtain animation results with realistic expressions and synchronized mouth movements. Benefiting from VOCASET Datset, VOCA [11] proposes a speaker-dependent facial animation pipeline that enables generalization across subjects. FaceFormer [4] introduces transformer [36] to capture long-term audio context and achieve temporary consistent animation performance. However, both of these two works lack attention to upper-face motions. To alleviate this problem, Meshtalk [18] learns a categorical latent space via a dedicated cross-modality loss. Subsequently, numerous works [5, 19, 12, 14, 20, 48, 49, 21, 22, 6] are proposed to further improve the realism. Codetalker [5] utilizes a discrete motion prior based on VQ-VAE [50] to avoid over-smoothed animation results. FaceDiffuser [14] and DiffPoseTalk [21] introduce diffusion model [51] to achieve diverse and realistic 3D facial motion generation, while DiffPoseTalk further considers the head pose generation. Meanwhile, Emotalk [19], EMOTE [20] and ExpCLIP [48] explore the emotional speech-driven 3D facial animation. However, all these methods focus on talking and are trained on talking datasets, leading to poor performance on the singing task. To this end, we collect the SingingHead dataset and benchmark existing 3D facial animation methods using the 3D facial motion of the dataset. An efficient baseline method that utilizes the background music is also proposed and benchmarked. Recently, Media2Face [22] shows some singing ability, however, it is not specialized for the singing task and is not open source. In contrast, we first propose the 4D singing head animation dataset and explore the 3D singing head animation task [52].
II-B Audio-driven 2D Portrait Animation
@@
2D portrait animation aims to generate lifelike videos that synchronized with the input audio [53, 54, 55, 56, 25, 15, 57, 1, 2, 16, 28, 58, 59, 17, 60, 61, 62]. Through the integration of a pre-trained Lip-Sync expert [63], Wav2Lip [24] enhances synchronization accuracy of the generated results. MekeItTalk [25] improves the realism via a speaker-aware audio encoder that enables personalized head pose movement. To enable natural head pose movement, PC-AVS [26] proposes a low-dimensional pose code to disentangle and control head poses. With the development in generation and reconstruction tasks, many methods [27, 28, 64, 65, 66, 67] introduce Neural Radiance fields [68] or diffusion model [51] to refine the quality of the generated videos. However, these methods typically suffer from less efficient generation processes due to the time-consuming volume rendering or diffusion sampling procedures. MODA [2] proposes a dual-attention module to estimate dense facial landmarks of deterministic and probabilistic movements to generate natural and detailed results. In recent years, some works [69, 16, 70] have explored the emotional talking head animation to achieve more realistic results. However, these methods mainly focus on speech audio, while another common type of audio, singing, remains under-explored. Since singing audio has a different rhythm compared to speech audio, and the expressions during singing are more exaggerated, it is challenging for current methods to directly generalize to the singing task. Therefore, we propose a large-scale singing head dataset for the development of singing head animation. Although some latest works [60, 61, 62] show some singing demos, they do not delve into the singing problem in depth. We are the first to study the singing head animation task based on our singing-specific dataset [52].
Figure 2: Comparison between talking and singing. (a) and (b) are the bar charts of the dominant frequency distribution and BPM distribution of audio, respectively. Our singing dataset has a wider range of dominant frequency and BPM distribution compared to talking datasets. (c) shows the chromagrams of a talking audio and a singing audio, which are very different and show that singing has more regular and stable pitch patterns as well as organized energy distribution. (d) is the distribution of mouth opening size calculated from facial landmarks, which shows that the mouth size is typically larger when singing compared to talking. All these demonstrate that singing is quite different from talking, and therefore needs to be studied separately.
II-C 2D and 3D Facial Animation Datasets
@@
Facial Animation Datasets play an important role in facial animation research, which can be divided into 2D and 3D datasets. 2D datasets [34, 9, 24, 71, 72, 73, 8] focus on collecting more diverse and higher quality photorealistic data. LRW dataset [24] contains 173 hours of low-resolution videos. To enlarge the data scale, the VoxCeleb dataset [71] collected more than 2k hours of videos from YouTube. In addition, several works [72, 34, 9] explore collecting emotional talking head data with emotion labels. The SAVEE [73] dataset is the first emotional dataset with only 4 actors. MEAD [9] dataset builds a high-resolution dataset with various video clips and viewpoints, covering a large number of emotional audio-visual corpus. Numerous 3D datasets [11, 18, 10, 7] have also been proposed for audio-driven 3D facial animation. MeshTalk dataset [18] contains detailed meshes topology and proposes a convolution neural network-based method. VOCASET dataset [11] constructs dataset by registering meshes to the FLAME [35] topology. However, these datasets are scale-limited. To this end, MMFace4D [7] collects a larger dataset consisting of 36 hours of data from 431 subjects. Recently, some works [19, 20, 21] have built 3D facial animation datasets from 2D portrait videos using 3D face reconstruction technologies to scale up the 3D dataset. However, all these datasets are collected for general talking and lack the generalization ability for singing tasks. To this end, a few works [32, 33] try to solve the singing task by collecting singing-specific datasets. Song2Face dataset [32] collects a 2-hour dataset with background music and 3D blendshapes from 7 singers. MusicFace [33] constructs a 3D dataset with only 6 singers by collecting the audio and intermediate videos separately. However, the existing two singing datasets are either small in size or of low quality, and neither provides 2D portrait singing videos, which limits their applications. Different from them, our SingingHead dataset is the first singing head dataset containing both 2D portrait videos and accurate 3D facial motions, with the greatest diversity (76 singers) and large scale (27 hours).
III SingingHead Dataset
Figure 3: Data collection pipeline. We use microphone, laptop, Azure Kinect camera, and light stage to collect raw data including singing audio, original songs downloaded from the Internet, singing video, and 3D head scan. Then, we process the raw data to obtain the cleaned singing audio, background music, portrait video, and 3D facial motion, which make up our SingingHead dataset.
Considering the lack of high-quality facial animation datasets specifically focused on singing, we collect SingingHead, a high-quality singing head animation dataset that contains both portrait video and accurate 3D facial motion with synchronized singing audio and background music. The dataset collection pipeline is shown in Fig. 3.
III-A Data Acquisition
@@
Design policy. We design the collection of SingingHead dataset by considering the following three points. (1) Variety of songs. For the diversity of the singing dataset, we carefully selected 8 common music categories, i.e., art music, country music, hip-hop, rock music, folk music, rhythm and blues, theater and drama, march, and asked the volunteers to choose 5 to 10 songs they were familiar with and fall to these categories. The pie chart in Fig. 4 (a) illustrates the categories and distributions of the songs we have finally collected. (2) Professionalism of volunteers. Different from talking, singing is a form of art that requires practice, which means that there is a great difference in singing skills between those who practice singing frequently and those who sing less. To this end, we select 50 professional singers and 26 amateurs as our volunteers, and recorded a total of 265 songs performed by professional volunteers and 182 songs performed by amateur volunteers, as illustrated in Fig. 4 (b). (3) Acquisition of dry sound. In normal circumstances, singers need to listen to the accompaniment or the original song when singing. In order to record the dry sound of singing, we equipped the volunteers with small Bluetooth earphones to play the accompaniment, and then recorded their voices via another microphone.
@@
Figure 4: Data Statistics. We show the pie charts of (a) the music category distribution, (b) the professional distribution of volunteers, (c) the gender distribution of volunteers, and (d) the language distribution, respectively.
III-B Data Processing
@@
The raw data includes RGB videos, singing dry sounds, original songs, and 3D scans. We processed the raw data as illustrated in Fig. 3 to get our final dataset SinginHead, which consists of singing audio, background music, portrait video, and 3D facial motion sequence.
@@
Video cropping and serialization. Since the video we collected has a wide field of view, including the entire upper body, and the volunteers routinely wiggled as they recorded the song, it is necessary to crop out the facial region and serialize the long recording to ensure the face remains in a stable bounding box. To this end, we first detect the bounding box of the facial region using MTCNN [75], then leave a margin around the bounding box to determine the crop box and crop the original video into square portrait video, finally, the portrait video is resized to $1024\times 1024$ as shown in Fig. 5. During the cropping process, we use Dlib [76] to detect the 68 landmarks every 6 frames, and calculate whether the landmarks on the edge are too close to the bounding box, if so, we cut the video at this frame, and start a new video sequence with a newly detected bounding box.
@@
3D facial motion reconstruction. We utilize the cropped portrait videos and the corresponding 3D scans to obtain accurate 3D singing facial motion sequences. Specifically, the 3D facial motion is represented using the expression and pose parameters of the commonly used parametric head model FLAME [35]. To estimate accurate FLAME parameters of the singing head, we first estimate the shape, expression, pose, and camera parameters separately for each frame in the portrait video with a SOTA monocular 3D face reconstruction method EMOCA v2 [77, 78, 79], and average the estimated shape parameters of all frames of a person as his/her shape parameter. Considering that estimating 3D shape from the 2D image is an ill-posed problem, and it is hard to obtain very accurate results, we further use the collected fine-gained 3D scans to fit a more accurate FLAME model by minimizing the 3D landmark loss and mesh-to-scan distance with the shape parameter initialized as the monocular estimated shape. Considering the FLAME parameters are estimated frame-by-frame, we utilize Box Plot to detect the outliers and replace them with the values of neighboring points to avoid discontinuous or incorrect motions. Concretely, we first calculate the average value of the estimated parameters over time, and then calculate Euclidean distance of each value to the corresponding average value and replace the outliers whose distances are outside the interval $[Q_{1}-s(Q_{3}-Q_{1}),Q_{3}+s(Q_{3}-Q_{1})]$, where $Q_{1}$ and $Q_{3}$ are the first and third quartile respectively, $s$ is the scale factor. Then, we smooth the detected 68 landmarks for the next optimization process by minimizing $||(\boldsymbol{lmk}_{2:T}-\boldsymbol{lmk}_{1:T-1})||_{2}^{2}$ using Adam optimizer [80]. Finally, we use the fitted FLAME shape as the final shape parameter, and reproject the estimated FLAME head back to the original portrait video to optimize the global orientation, neck pose, and camera parameters by minimizing:
$\mathcal{L}=\lambda_{l}\mathcal{L}_{lmk}+\lambda_{s}\mathcal{L}_{s}+\lambda_{r}\mathcal{L}_{reg},$ (1)
where $\mathcal{L}_{lmk}$, $\mathcal{L}_{s}$ and $\mathcal{L}_{reg}$ denotes landmark reprojection loss, smoothing loss, and regularization term, respectively. $\lambda_{l}$, $\lambda_{s}$ and $\lambda_{r}$ are the weight coefficients. Specifically, landmark reprojection loss $\mathcal{L}_{lmk}$ is calculated by:
$\mathcal{L}_{lmk}=\left\|(\boldsymbol{\hat{lmk}}-\boldsymbol{lmk})\right\|_{2}^{2},$ (2)
where $\boldsymbol{\hat{lmk}}$ and $\boldsymbol{lmk}$ represent ground truth and reprojected 2D landmarks respectively. Smoothing loss $\mathcal{L}_{s}$ is the sum of the smoothing losses of global orientation, neck pose, camera and bounding box parameters with same formulation: $||(\boldsymbol{X}_{2:T}-\boldsymbol{X}_{1:T-1})||_{2}^{2}$ where $\boldsymbol{X}$ represent any one of these variables. The regularization term $\mathcal{L}_{reg}=||\boldsymbol{\theta_{g}}||_{2}^{2}+||\boldsymbol{\theta_{n}}||_{2}^{2}$, where $\boldsymbol{\theta_{g}}\in\mathbb{R}^{3}$ represents the global orientation and $\boldsymbol{\theta_{n}}\in\mathbb{R}^{3}$ is the FLAME neck pose parameter.
After these steps, we can obtain accurate and smooth 3D singing facial motions represented by 59-dimensional FLAME [35] parameters (i.e., 3 dimensions for global orientation, 50 dimensions for expression, 3 dimensions for jaw pose and 3 dimensions for neck pose). Besides, all the subjects have their 100-dimensional FLAME shape parameter. Some key frames of our 3D facial motions are shown in the second row of Fig. 5, demonstrating the high precision of the reconstructed 3D facial motions.
Figure 5: Visualization of SingingHead dataset. We show some key frames from singing portrait videos and the corresponding 3D facial motions.
III-C Dataset Analysis
Data statistics. According to the aforementioned pipeline, we record a total of 448 raw video sequences with synchronized singing audio from 76 subjects. Each sequence comes from a complete song, ranging from 27 to 457 seconds. Fig. 4 shows pie charts of the music category distribution and language distribution of these recordings, as well as the distribution of whether the subjects are professional and their gender. Then, portrait videos are cut into smaller segments to keep the facial region within the cropped bounding box, resulting in 10223 video sequences. The singing audio, background music, and 3D facial motion are then segmented according to the video segments. Sequences less than 3s are all removed. Therefore, the SingingHead dataset consists of 1952 synchronized singing videos, 3D facial motions, singing audios, and background music ranging from 3 to 432 seconds. For convenience, we also cut these sequences into equal-length 8s segments for benchmarking and training our method, resulting in a total of 12196 8s sequences.
Differences from talking head datasets. As shown in Fig. 2, our singing dataset is quite different from existing talking datasets both in terms of audio features and facial movements, so it is necessary to collect the singing-specific dataset. To be specific, singing audio has a wider range of frequency and BPM (Beats Per Minute) due to the diversity of songs and their broader range of tones and rhythms. In talking datasets, people usually speak at moderate speaking rate and intonation, so that the frequencies and BPMs of talking audio are concentrated in the lower frequency range and middle range, respectively. Different from talking datasets, we ask the volunteers to sing a variety of songs including fast songs, slow songs, bass songs, and treble songs, which results in a much wider range of frequency and BPM distribution of the singing audio in our SingingHead dataset. Moreover, people often exhibit more exaggerated expressions when singing compared to talking for artistic performance and emotional communication. For instance, the mouth opening size is typically larger when singing compared to talking as shown in Fig. 2 (d). Besides, there tend to be more head movements when singing than when talking. Therefore, it is essential for models to learn these movements from singing-specific datasets to achieve vivid and realistic singing head animation.
IV Method